OpenStudy (anonymous):

The cost for Andrea to produce several units of a certain item is $336 plus $12 per unit. She sells this item for $28 per unit. How many units of the item must she produce and sell so that the amount of money she earns in sales equals the amount of money she spends in production costs? How much money would that be? She would have to produce and sell ______ units of the item. Her costs and her earnings would both equal $___.

OpenStudy (anonymous):

@perl

OpenStudy (anonymous):

@gorv

OpenStudy (anonymous):

I really need some help :(

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

@LunyMoony

OpenStudy (anonymous):

Thank you for coming :)

OpenStudy (anonymous):

I really need help

OpenStudy (lunymoony):

Ok

OpenStudy (anonymous):

do you have any ideas?

OpenStudy (anonymous):

You there?

OpenStudy (lunymoony):

yes

OpenStudy (anonymous):

.........

OpenStudy (lunymoony):

she will have to sell 21 items to cover her costs y = several units, x = the certain item (y)x, the certain item times its amount the several units of a certain item = $336, + $12(y) (y)x = $336 +$12(y) She sells the item for $28 each, x = $28 (y)28 = 336 + 12(y) 28y = 336 + 12y take -12y from each side 28y - 12y = 336 16y = 336 y = 21

OpenStudy (lunymoony):

@Ri0T_L1on

OpenStudy (anonymous):

ok so she would have to produce 21 items each?

OpenStudy (anonymous):

and then what would the earnings be?

OpenStudy (anonymous):

Thank you soooo much i got it :) you multiply 21 x 28 and you get your answer

OpenStudy (anonymous):

and i gave you a medal:)

OpenStudy (lunymoony):

Sry, had to reload my page before I realized you texted me back, awesome, glad you got it

OpenStudy (anonymous):

bye :)